Carol, a Grade 3 learner, has a heart rate of 56 beats per minute. Calculate how many times her heart will beat in 5 minutes, 30 seconds, 1 hour?
Answer by Theo(13342)

60 seconds = 1 minute.
30 seconds = 1/2 * 60 seconds = 1/2 of a minute.
1 hour = 60 minutes.

carol has a heart rate of 56 beats per minute.

in 5 minutes, her heart will beat 5 * 56 = 280 times.

in 30 seconds, her heart will beat 1/2 * 56 = 28 times.

in 1 hour, her heart will beat 60 * 56 = 3360 times.